OpenStudy (anonymous):

Calculate the amount of zirconium produced (in kg) if a reactor were charged with 310.0 kg ZrCl4 and 29.0 kg Na.

OpenStudy (anonymous):

Similar question: Calculate the amount of zirconium produced (in kg) if a reactor were charged with 491.0 kg ZrCl4 and 49.0 kg Na. ZrCl4+4Na→4NaCl+Zr mw (ZrCl4) =91.22 + 4(35.45) =233.02 gm/mole mw (Na) = 22.44 gm/mole 491.0 x〖10〗^23 gm / 233.02 gm/mole = 2107.11 mole ZrCl4 49.0 x 〖10〗^23 gm / 22.44 gm/mole = 2183.60 mole Na =>Require 2183.60/4 = 545.75 mole limiting reagent ZrCl4 545.75 x 91.22 = 49783.315 gm Zr = 49.78 kg Zr.